# Technologies for Modulation of Visible and Near-UV Light and their Applications

### v1

@@ III.5 Summary
In contrast to the devices discussed in Sec. II, the integrated photonic systems discussed in this section are less mature, but promise scalability due to significant improvements in size and per-channel power. Fig. 4(a) presents a qualitative overview of the four main integrated technologies discussed in this review. There is a clear trade-off between different figures of merit, most obviously between the control power and speed of modulation. Similarly, Fig. 4(b) presents an overview of all technologies and publications discussed in this review with respect to operating wavelength and the modulation speed. While there has been significant work in developing visible light integrated photonics, there is a clear gap in demonstrating fast, low-wavelength modulation, particularly critical for quantum applications. Moreover, most published works focus on single-device characteristics and systems-scale optimization will be needed before they become ubiquitous in application spaces.
@@ IV Outlook and Conclusion
Indeed, hybrid solutions are already being employed to provide simultaneously high temporal and spatial bandwidth. This review has mainly focused on single-beam modulation. However, for many neutral-atom quantum computing, biological control, and augmented-reality display applications, the spatial bandwidth, or the number of individually controllable channels, is also important. There is currently no technology enabling both high-speed modulation and high-resolution spatial modulation. AOMs, EOMs, and integrated photonics allow for sub-microsecond modulation but are limited to a few channels. The SLMs based on MEMS or LCoS dynamic surfaces enable light modulation in the spatial domain with millions of resolvable spots, but their modulation speed is limited to tens of kilohertz. This phenomenological trade-off has inspired research on developing high-speed SLMs to achieve full spatial and temportal control of incident light with resonant structures and metasurfaces [241, 242].
Additionally, high-speed single or few-channel modulators integrated with high-resolution SLMs offer an intermediate solution. Specifically, an angle- (or position-) multiplexed hologram displayed on a high-resolution SLM is illuminated by different spatial modes whose amplitude and phase are modulated by an array of high-speed optical modulators [243]. These intermediate solutions are used for high-speed optical mode sorting [244], neutral-atom quantum information processing [243], and the individual beam shaping of modulated beam arrays [186].
Finally, the results presented in Sec. III have, for the most part, been device-scale demonstrations. However, all of the applications discussed in Sec. I necessitate multiple channels and full system packaging. This includes coupling light out of the chip with high efficiency, simultaneous control of multiple channels with low crosstalk, and integrated driving electronics. Additionally, while we focused on five figures of merit in this review, it will also be important to improve other capabilities such as on-off isolation, optical power, and device crosstalk. As these technologies mature from device-scale demonstrations to adoption by the applications community it will be necessary to optimize system-scale specifications.

### v2

@@ III.6 Summary
In contrast to the devices discussed in Sec. II, the integrated photonic systems discussed in this section are less mature, but promise scalability due to significant improvements in size and per-channel power. Fig. 9(a) presents a qualitative overview of the four main integrated technologies discussed in this review. There is a clear trade-off between different figures of merit, most obviously between the control power and speed of modulation.
Similarly, Fig. 9(b) presents an overview of all technologies and publications discussed in this review with respect to operating wavelength and the modulation speed. While there has been significant work in developing visible light integrated photonics, there is a clear gap in demonstrating fast, low-wavelength modulation, particularly critical for quantum applications. Moreover, most published works focus on single-device characteristics and systems-scale optimization will be needed before they become ubiquitous in application spaces.
@@ IV Outlook and Conclusion
Indeed, hybrid solutions are already being employed to provide simultaneously high temporal and spatial bandwidth. For many neutral-atom quantum computing, biological control, and augmented-reality display applications, the spatial bandwidth, or the number of individually controllable channels, is also important. There is currently no technology enabling both high-speed modulation and high-resolution spatial modulation. AOMs, EOMs, and integrated photonics allow for sub-microsecond modulation but are limited to a few channels. The SLMs based on MEMS or LCoS dynamic surfaces enable light modulation in the spatial domain with millions of resolvable spots, but their modulation speed is limited to tens of kilohertz.
High-speed single or few-channel modulators integrated with high-resolution SLMs offer an intermediate solution. Specifically, an angle- (or position-) multiplexed hologram displayed on a high-resolution SLM can be illuminated by different spatial modes whose amplitude and phase are modulated by an array of high-speed optical modulators [252]. These intermediate solutions are used for high-speed optical mode sorting [253], neutral-atom quantum information processing [252], and the individual beam shaping of modulated beam arrays [195].
Finally, the results presented in Sec. III have, for the most part, been device-scale demonstrations. However, all of the applications discussed in Sec. I necessitate multiple channels and full system packaging. This includes coupling light out of the chip with high efficiency, simultaneous control of multiple channels with low crosstalk, and integrated driving electronics. Additionally, while we focused on five figures of merit in this review, it will also be important to improve other capabilities such as on-off isolation, optical power handling, and device crosstalk. As these technologies mature from device-scale demonstrations to adoption by the applications community it will be necessary to optimize system-scale specifications.